Use the method of cylinders to determine the volume of the solid obtained by rotating the region bounded by x=(y-2)^2, the x-axis and the y-axis about the x-axis.


1
Expert's answer
2021-04-13T15:45:22-0400

We can calculate the are by using method of cylinder

"A(y)=2\pi\times y(y-2)^2\\"

We can integrate this from 0 to 2 because these are common points

"V(y)" "= \int _0^22\pi\times y(y-2)^2dy"


"\Rightarrow [ 157
\dfrac{(
y
−
2
)
^3
(
3
y
+
2
)}
{300}]_0^2"


"\Rightarrow 8.3733"